# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 13 Wigner collaborates with Deutsch
Based on the framework developed in Section 7, we can analyze more complicated situations than considered in the previous sections. In this section, as such a example, we consider the combination of Wigner’s friend and Deutsch’s thought experiment. The system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ participate in the experiment. Each of three observers makes their measurements as before but in the order of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$, where the observer $\mathcal{D}$ do nothing to the observer $\mathcal{W}$. We call this combination the Wigner-Deutsch collaboration. In this section, we describe the Wigner-Deutsch collaboration in detail in the terminology of the conventional quantum mechanics from scratch.
@@
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
@@
We define the state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
After the measurement $\mathcal{M}^{\mathcal{W}}$ by $\mathcal{W}$ on the observer $\mathcal{F}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
@@
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{W}}$, we see that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome 0 with certainly in the case of the occurrence of the possibility (i), and it gives the outcome $1$ with certainly in the case of the occurrence of the possibility (ii). The measurement $\mathcal{M}^{\mathcal{W}}$ does not change the state of the composite system $\mathcal{S}+\mathcal{F}$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. In this Case 1, we can consider the joint probability $P_{\mathrm{c1}}(k,l,m)$ that the measurement $\mathcal{M}^{\mathcal{F}}$ gives the outcome $k$ and then the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $k=0,1$, $l=0,1,2$, and $m=+,$-. We have that $P_{\mathrm{c1}}(k,l,m)=\delta_{k,l}/4$ for every $k=0,1$, $l=0,1,2$, and $m=+,$-. Then the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
$|\Psi^{\mathcal{S}+\mathcal{F}}[-]\rangle:=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle-|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$
@@
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{W}}$. Thus, in this Case 2, we can consider the joint probability $P_{\mathrm{c2}}(l,m)$ that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $l=0,1,2$ and $m=+,$-. We have that $P_{\mathrm{c2}}(l,m)=1/4$ for every $l=0,1$ and $m=+,$-, and $P_{\mathrm{c2}}(2,m)=0$ for every $m=+,$-. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2. In the same manner as in Case 1, the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
Thus, if we ignore the measurement results of the measurement $\mathcal{M}^{\mathcal{F}}$ in Case 1, there is no difference between Case 1 and Case 2 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$. Actually, we have that
@@ 14 Analysis of the Wigner-Deutsch collaboration based on the principle of typicality
In this section, we make an analysis of the Wigner-Deutsch collaboration described in the preceding section, in our framework of quantum mechanics based on Postulates 5, 6, and 7, together with Postulates 1, 2, and 3.
@@
where $I_{\mathcal{S}}$ and $I_{\mathcal{W}}$ denote the identity operators acting on $\mathcal{H}_{\mathcal{S}}$ and $\mathcal{H}_{\mathcal{W}}$, respectively. Thus, it follows from (62) that
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. We denote the set $\{0,1\}\times\{0,1,2\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
@@
Thus, the finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ forms measurement operators. Then, using (66) we see that, as a whole, the sequential applications of $U_{\mathrm{I}}$, $U_{\mathrm{II}}$, and $U_{\mathrm{III}}$ result in:
@@ 14.1 Application of the principle of typicality
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$, $\beta$, and $\gamma$ to denote the infinite sequences over $\{0,1\}$, $\{0,1,2\}$, and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (71) and Corollary 4.4 that
@@
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the equality follows from (70). Since $\omega$ is Martin-Löf $P$-random, it follows from (72) and Theorem 4.2 that for every $(k,l,m)\in\Omega_{c}$ it holds that
@@
where $N_{k,l,m}(\sigma)$ denotes the number of the occurrences of $(k,l,m)$ in $\sigma$ for every $(k,l,m)\in\Omega_{c}$ and every $\sigma\in\Omega_{c}^{*}$. Thus, in our world $\omega$, the following holds for each $(k,l,m)\in\Omega_{c}$: In one-eighth of the infinite repetitions of the experiment, the state of $\mathcal{S}$ is $|k\rangle$ and the observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ record the values $k$, $l$, and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that this situation is different from both Case 1 and Case 2 in Section 13, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given, in particular, in the Born rule, Postulate 4 (i).
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ is confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (73), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is
@@
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$, $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$, $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. On the one hand, applying $(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}})$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (63) and (55), respectively. On the other hand, the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. But it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (73) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
@@
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
@@
Now, we can apply Postulate 7 recursively, based on (79). Namely, we can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Recall that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ form PVMs. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments where the observer $\mathcal{D}$ is ignored results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (56) and (55), respectively. On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (79) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{W}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$P\left(|\Phi^{\mathcal{F}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\beta(n)]\rangle\langle\Phi^{\mathcal{F}}[\beta(n)]\rvert\otimes I_{\mathcal{W}}.$
@@
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Hence, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the observer $\mathcal{F}$ surely records the value $\beta(n)$ regarding the system $\mathcal{S}$ whose state is $|\beta(n)\rangle$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in the $n$th repetition of the experiment, in accordance with our common sense. Note that we can also derive the result (82) from (79) using Postulate 6 in the same manner as we derived the result (35) in Section 9.1, instead of using Postulate 7.
@@
We can then see that the this situation regarding $\mathcal{M}^{\mathcal{F}}$ and $\mathcal{M}^{\mathcal{W}}$ coincides with Case 1 in Section 13, although the statement of Case 1 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.

### v2

@@ 13 Wigner collaborates with Deutsch
Based on the framework developed in Section 7, we can analyze more complicated situations than considered in the previous sections. In this section, as such an example, we consider a combination of Wigner’s friend and Deutsch’s thought experiment. The system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ participate in the experiment. Each of three observers makes their measurements as before but in the order of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$, where the observer $\mathcal{D}$ do nothing to the observer $\mathcal{W}$. We call this combination the Wigner-Deutsch collaboration. In this section, we describe the Wigner-Deutsch collaboration in detail in the terminology of the conventional quantum mechanics from scratch.
@@
After the measurement $\mathcal{M}^{\mathcal{F}}$ performed by the observer $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
@@
We define a state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
After the measurement $\mathcal{M}^{\mathcal{W}}$ performed by the observer $\mathcal{W}$ on the observer $\mathcal{F}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs immediately after the measurement $\mathcal{M}^{\mathcal{F}}$:
@@
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{W}}$, we see that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome 0 with certainty in the case of the occurrence of the possibility (i), and it gives the outcome $1$ with certainty in the case of the occurrence of the possibility (ii). The measurement $\mathcal{M}^{\mathcal{W}}$ does not change the state of the composite system $\mathcal{S}+\mathcal{F}$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. In this Case 1, we can consider the joint probability $P_{\mathrm{c1}}(k,l,m)$ that the measurement $\mathcal{M}^{\mathcal{F}}$ gives the outcome $k$ and then the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $k=0,1$, $l=0,1,2$, and $m=+,$-. We have that $P_{\mathrm{c1}}(k,l,m)=\delta_{k,l}/4$ for every $k=0,1$, $l=0,1,2$, and $m=+,$-. Then, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
$|\Psi^{\mathcal{S}+\mathcal{F}}[-]\rangle:=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle-|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$ (62)
@@
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{W}}$. Thus, in this Case 2, we can consider the joint probability $P_{\mathrm{c2}}(l,m)$ that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $l=0,1,2$ and $m=+,$-. We have that $P_{\mathrm{c2}}(l,m)=1/4$ for every $l=0,1$ and $m=+,$-, and $P_{\mathrm{c2}}(2,m)=0$ for every $m=+,$-. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2. In the same manner as in Case 1, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
Thus, if we ignore the measurement results of the measurement $\mathcal{M}^{\mathcal{F}}$ in Case 1, there is no difference between Case 1 and Case 2 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$. Actually, we can see that
@@ 14 Analysis of the Wigner-Deutsch collaboration based on the principle of typicality
In this section, we make an analysis of the Wigner-Deutsch collaboration, which is described in the preceding section, in our framework of quantum mechanics based on Postulates 5, 6, and 7, together with Postulates 1, 2, and 3.
@@
where $I_{\mathcal{S}}$ and $I_{\mathcal{W}}$ denote the identity operators acting on $\mathcal{H}_{\mathcal{S}}$ and $\mathcal{H}_{\mathcal{W}}$, respectively. Thus, it follows from (59), (62), and (64) that
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. Actually, using (66) and (67) we can check that (68) holds certainly in the case where $|\Psi\rangle=|0\rangle$ or $|\Psi\rangle=|1\rangle$. Thus, due to linearity we have that (68) holds for an arbitrary $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, as desired.
We denote the set $\{0,1\}\times\{0,1,2\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
@@
Thus, the finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ satisfies the completeness equation, and therefore forms measurement operators. Then, using (68) we see that, as a whole, the sequential applications of $U_{\mathrm{I}}$, $U_{\mathrm{II}}$, and $U_{\mathrm{III}}$ result in:
@@ 14.1 Application of the principle of typicality
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiment) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ which is being generated by the infinite repetition of the measurement described by the measurement operators $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$, $\beta$, and $\gamma$ to denote the infinite sequences over $\{0,1\}$, $\{0,1,2\}$, and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then, since $\omega$ is Martin-Löf $P$-random, it follows from (73) and Corollary 4.4 that
@@
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. Thus, it follows from Postulate 5 and (70) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the equality follows from (72). Since $\omega$ is Martin-Löf $P$-random, it follows from Theorem 4.2 and (74) that for every $(k,l,m)\in\Omega_{c}$ it holds that
@@
where $N_{k,l,m}(\omega\!\!\restriction_{n})$ denotes the number of the occurrences of $(k,l,m)$ in the prefix of $\omega$ of length $n$. Thus, in our world $\omega$, the following holds for each $(k,l,m)\in\Omega_{c}$: In one-eighth of the infinite repetitions of the experiment, the state of $\mathcal{S}$ is $|k\rangle$ and the observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ record the values $k$, $l$, and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that this situation is different from both Case 1 and Case 2 in Section 13, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. According to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ are unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ are confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (75), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is
@@
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that each of the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$, $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$, and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ forms a PVM. On the one hand, applying $(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}})$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (65) and (56), respectively. On the other hand, the final states of the observer $\mathcal{W}$ are unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. But it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$. Here “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (75) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
@@
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
@@
Now, we can apply Postulate 7 recursively, based on (81). Namely, we can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Recall that each of the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ forms a PVM. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ in the repeated once of the experiments where the observer $\mathcal{D}$ is ignored results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (57) and (56), respectively. On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ are unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ are confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (81) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{W}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$P\left(|\Phi^{\mathcal{F}}[\beta(n)]\rangle\right)=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\beta(n)]\rangle\langle\Phi^{\mathcal{F}}[\beta(n)]\rvert\otimes I_{\mathcal{W}}.$
@@
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Hence, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the system $\mathcal{S}$ is in the state $|\beta(n)\rangle$ and the observer $\mathcal{F}$ surely records the value $\beta(n)$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in the $n$th repetition of the experiment. Note that we can also derive the result (84) from (81) using Postulate 6 in the same manner as we derived the result (35) in Section 9.1, instead of using Postulate 7.
@@
We can then see that this situation regarding $\mathcal{M}^{\mathcal{F}}$ and $\mathcal{M}^{\mathcal{W}}$ coincides with Case 1 in Section 13, although the statement of Case 1 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.